OpenStudy (anonymous):

When a flare is fired upward at 58.8 m/s, its height, h meters, is given by the equation h=-4.9t + 58.8t, where t seconds is the time since firing. a)determine the max height of the flare and the time it takes to reach this height. b) For how many seconds is the flare higher than 98m?

OpenStudy (anonymous):

There's something wrong with your equation. It's probably supposed to be: -4.9t^2

OpenStudy (anonymous):

a) you need to find the vertex which = -b/2a. that will be the time at which it hits the max

OpenStudy (anonymous):

@ SmoothMath yah, sorry.. It's supposed to be squared. :p

OpenStudy (anonymous):

Well then. Since the height is -4.9t^2 +58.8t, figure out when the height is 98 by solving the equation -4.9t^2 + 58.8t = 98 There should be two solutions to this equation, one is the time it passes that height going UP, and the other is the time that it passes that height going DOWN.

OpenStudy (anonymous):

\(h(t)=-4.9t^2 + 58.8t\) \(-\frac{b}{2a}=-\frac{58.8}{2\times (-4.9)}=6\) is the value that will give the max

OpenStudy (anonymous):

To solve that equation, you should probably use the quadratic formula. Simply rewrite in the form Ax^2 + Bx + C = 0, and then plug A, B, and C into the quadratic formula =D

OpenStudy (anonymous):

As for the max, Satellite just gave you the answer, so that's cool.

OpenStudy (anonymous):

second one solve \[-4.9t^2 + 58.8t \geq 98\] \[-4.9t^2 + 58.8t -98\geq 0\] or better yes \[4.9t^2-58.8+98\leq 0\]

OpenStudy (anonymous):

no in fact 6 is not the answer, 6 is the value of \(t\) that will give you the answer

OpenStudy (anonymous):

and for last one, do not use the quadratic formula, it will give you a big mess. the problem has been cooked to be easy. step one is divide everything by \(4.9\) and then you will have a simple quadratic to solve by factoring

OpenStudy (anonymous):

Because it's so obvious to the casual algebra student that this equation is divisible by 4.9. -_-

OpenStudy (anonymous):

Thanks guys c: I got 529.2?